A box with a total surface area of 0.55 m2
and
a wall thickness of 4.51 cm is made of an insulating material. A 16 W electric heater inside
the box maintains the inside temperature at
7.5
◦C above the outside temperature.
Find the thermal conductivity of the insulating material.
Answer in units of W/m ·
◦ C.
1
Expert's answer
2020-04-29T09:48:57-0400
"P=kA\frac{\Delta T}{h}"

"16=(0.55)\frac{7.5}{0.0451}k"

"k=0.175\frac{W}{m\degree C}"